OpenStudy (emac):

Suppose that x=x(t) and y=y(t) are both functions of t. If y=2x^2-1 and dx/dt=6 when x=11 what is dy/dt?

OpenStudy (steve816):

Step 1: Can you take the derivative of the equation y=2x^2-1

OpenStudy (emac):

y'=4x

OpenStudy (irishboy123):

\(\left. \dfrac{dy}{dt} \right|_{x = 11} = \left. \dfrac{dy}{dx} \right|_{x = 11} \cdot \left. \dfrac{dx}{dt}\right|_{x = 11}\) \(= \left. 4x \right|_{x = 11} \cdot \left. \dfrac{dx}{dt}\right|_{x = 11}\) \(= 44 \cdot 6\)

OpenStudy (emac):

Thank you!! that makes sense I was having a hard time interpreting the instructions